Volume of co2 obtained at stp by the complete decomposition of 9.85 g na2co3 is.

Grade:11

8 Answers

Ashutosh Sharma
8 years ago

Na2CO3⇒ Na2O + CO2
1 mole of sodium carbonate give 1 mole of carbon dioxide.
as 105.98g sodium carbonate gives 44.01g carbon dioxide
therefore, 9.85 g sodium carbonate gives = 9.85 × 44.01 =4.09 g of CO2 105.98
Moles of CO2 mole = weight =4.09/44.01= 0.092 mole
Molar mass

1 mole of gas = 22.4 L gas
Therefore,0.092 mole of CO2 = 22.4 × 0.092 = 2.06 L CO2.

Rahul Ray
7 years ago
Equation:Na2co3➡️Na2o+co2
On calculating this reaction one might get answer 2.08 litres but as Na2co3 is soda ash and it does not decomposes on heating even to redness
And co2 does not evolved
  So the answer will be zero litres
Anand
6 years ago
Na2co3 requires temperature above 500 degree Celsius to decompose that is not possible at STP so the answer will be zero cow evolved
SJ Nihal
6 years ago
Na2CO3 on decomposition gives CO2 .This is unfair if , 9.85 gm of sodium carbonate decompose and 2.08 litres CO2 formed as per calculation. As it does not decompose formally by heating .
Prabal Sarkar
5 years ago
Na2CO3 require a high temperature(500℃).As the reaction happen in STP. So it can not get proper heat to make gasious CO2. So the volume of CO2 is Zero.....
Bashir
5 years ago
It's answer will be zero bcz there will be no effect on Na2CO3 on decomposition
It need lots of energy to break the bonds ofNa2CO3 which is not possible in ordinary heating.
Rishi Sharma
3 years ago
Hello Soumyadee,
The solution of the problem is as follows
Na2CO3⇒ Na2O + CO2
1 mole of sodium carbonate give 1 mole of carbon dioxide
as 106 g sodium carbonate gives 44g carbon dioxide
therefore, 9.85 g sodium carbonate gives = 9.85 × 44 / 106 = 4.09 g of CO2
Moles of CO2 mole =4.09/44 = 0.092 mole
Molar mass 1 mole of gas = 22.4 L gas
Therefore, 0.092 mole of CO2 = 22.4 × 0.092 = 2.06 L CO2.
Thanks
I hope above solution will clear your all doubts.
Please feel free to post and ask as much doubts as possible.
All the best.


Farhan
2 years ago
Answer is zero as na2co3 does not decompose at STP. A temperature of above 500°C is required for Na2CO3 to break bods. Moreover it is soda ash. 
 
Hope the answer will be useful.